The diameter of a wire is (0.56+_0.01)mm, what is the cross sectional area of the wire?
1
Expert's answer
2018-03-23T10:18:07-0400
A=(πd^2)/4=π/4 (0.56)^2=0.25 mm^2.
Δd/d=0.01/0.56=1/56
ΔA/A=2 Δd/d=2/56=1/28
ΔA=0.25/28=0.01 mm^2.
The cross sectional area of the wire is
(0.25±0.01) mm^2.